# Detecting Harmful Memes with Decoupled Understanding and Guided CoT Reasoning

### v1

@@
Abstract
Detecting harmful memes is essential for maintaining the integrity of online environments. However, current approaches often struggle with resource efficiency, flexibility, or explainability, limiting their practical deployment in content moderation systems. To address these challenges, we introduce U-CoT+, a novel framework for harmful meme detection. Instead of relying solely on prompting or fine-tuning multimodal models, we first develop a high-fidelity meme-to-text pipeline that converts visual memes into detail-preserving textual descriptions. This design decouples meme interpretation from meme classification, thus avoiding immediate reasoning over complex raw visual content and enabling resource-efficient harmful meme detection with general large language models (LLMs). Building on these textual descriptions, we further incorporate targeted, interpretable human-crafted guidelines to guide models’ reasoning under zero-shot CoT prompting. As such, this framework allows for easy adaptation to different harmfulness detection criteria across platforms, regions, and over time, offering high flexibility and explainability. Extensive experiments on seven benchmark datasets validate the effectiveness of our framework, highlighting its potential for explainable and low-resource harmful meme detection using small-scale LLMs.   [^1] Codes and data are available at: https://anonymous.4open.science/r/HMC-AF2B/README.md.
Detecting Harmful Memes with Decoupled Understanding and Guided CoT Reasoning
Fengjun Pan  Anh Tuan Luu thanks: Corresponding Authors.  Xiaobao Wu 11footnotemark: 1 Nanyang Technological University {fengjun001, anhtuan.luu, xiaobao002}@ntu.edu.sg
This paper includes content for demonstration purposes that some readers may find disturbing.
1 Introduction
@@
Memes have emerged to be a popular communication medium on social media platforms Nguyen and Ng (2024); Joshi et al. (2024). By combining images with overlaid text Dancygier and Vandelanotte (2017), they can convey rich information in a visually engaging and immediate manner. However, despite their often humorous nature, memes can be exploited to disseminate harmful content, such as hate speech Kiela et al. (2020); Davidson et al. (2019), toxic misinformation Lu et al. (2024); Attanasio et al. (2022), or inflammatory opinions Liu et al. (2022) etc. under the guise of humor. As such, harmful memes pose a threat to the safety and harmony of online environments Pandiani et al. (2024); Sharma et al. (2022), thus motivating the development of automated harmful meme detection methods, which typically follow supervised fine-tuning (SFT).
@@
However, the adoption of these methods in real-world content moderation systems is limited by three critical issues: (i) Low resource efficiency. Supervised approaches require large amounts of labeled training data to perform well Cao et al. (2023b); Burbi et al. (2023); Shah et al. (2024); Lin et al. (2024a); Cao et al. (2023a); Kumar and Nandakumar (2022); Pramanick et al. (2021b), but collecting such data is often time-consuming, costly, and potentially biased, as harmful memes with nuanced and implicit meanings can be difficult for human annotators to label consistently without being influenced by their individual knowledge and subjective perceptions. Besides, previous methods often rely on expensive advanced multimodal models Huang et al. (2024); Ji et al. (2024); Lin et al. (2024a) such as GPT-4Hurst et al. (2024), posing significant accessibility challenges for practitioners with limited resources. (i) Low flexibility. The classification criteria distinguishing harmful and harmless memes often vary across sociocultural contexts Pramanick et al. (2021a); Sharma et al. (2022) and differ significantly across platforms, regions, and over time. Once trained, supervised models becomes difficult to adapt to new criteria and out-of-domain meme types without undergoing costly data collection and retraining Mei et al. (2025); Cao et al. (2024). (i) Low explainability. Most existing supervised methods operate as black boxes, offering no human-interpretable reasoning behind their predictions Burkart and Huber (2021); Yang et al. (2024); Mei et al. (2023); Burbi et al. (2023). This lack of transparency in decision-making undermines user trust and the reliability for deployment in sensitive or regulated scenarios Sasikala and Sachan (2024), where rationales are essential to support subsequent verification and moderation feedback.
@@
To address these limitations, in this paper, we propose U-CoT+, an unimodal guided CoT-based zero-shot harmful meme detection framework. It converts the default multimodal classification setup back to unimodal text-based inference that enables general Large Language Models (LLMs) to perform zero-shot harmful meme detection solely based on natural language input. Specifically, we first propose a High-Fidelity Meme2Text pipeline. Instead of directly prompting Large Multimodal Models (LMMs) to make classification decisions as before, we query LMMs multiple times to extract fine-grained visual details from memes Cao et al. (2023a); Ji et al. (2024), including sensitive identity characteristics of human subjects e.g., race, gender, appearance, and disability status. We then leverage an LLM to integrate these visual details into a unified and coherent description of the meme’s content. As such, we effectively decouple meme understanding from the reasoning of meme classification. This reduces the task complexity by avoiding direct reasoning over complicated visual elements, thereby enabling zero-shot harmful meme detection using general-purpose LLMs.
Furthermore, we introduce a Guided CoT (Chain-of-Thoughts Wei et al. (2022)) prompting approach. Rather than relying solely on the inherent knowledge embedded in LLMs, we provide explicit, human-crafted guidelines that define the fine-grained, targeted criteria for detecting harmful content in memes. LLMs perform classification by reasoning over meme descriptions in conjunction with the classification guidelines. This encourages consistent, interpretable, and aligned classifications, and more importantly, enables the framework to easily and effectively adapt to varying harmful meme criteria and a wide range of memes from diverse sociocultural contexts. Through extensive experiments on benchmark datasets, we show that our proposed LLMs-based unimodal harmful meme detection framework under a zero-shot setting can achieve performance highly comparable to state-of-the-art full-shot supervised fine-tuned baselines. Notably, with guided CoT prompting, small-scale LLMs (such as Mistral-12B Mistral (2024) and Qwen2.5-14B Qwen et al. (2025)) perform competitively through zero-shot inference, matching and even outperforming GPT-4o and other low-resource baselines.
Figure 1: Illustration of previous methods and our U-CoT+. Previous methods follow either fully supervised or low-resource settings, fine-tuning LMMs/LLMs with labeled data or prompting advanced LMMs (e.g., GPT-4) with (or without) few-shot examples or retrieval-augmented mechanisms. These methods do not necessarily guarantee predictions that include explicit reasoning. U-CoT+ employs a High-fidelity Meme2Text pipeline to convert the multimodal harmful meme detection task into a unimodal, text-only setting, and further enhances LLM performance through Unimodal Guided CoT Prompting. An output example given by Qwen2.5-14B under U-CoT+ is presented in “Step-by-step Reasoning”.
We summarize our main contributions as three-fold:
@@
  • We propose a resource-efficient, unimodal harmful meme detection framework that effectively leverages the inherent reasoning abilities of pre-trained LLMs to enable interpretable meme classification in low-resource settings. By performing zero-shot inference on high-fidelity meme descriptions under guided CoT prompting, our framework can easily adapt to different harmful meme criteria across diverse contexts.
@@
  • We contribute a set of interpretable and targeted harmful meme detection guidelines that capture diverse features of harmful memes across different contexts. We demonstrate that these high-quality, human-crafted guidelines can significantly enhance the zero-shot performance of LLMs in harmful meme detection.
  • We conduct extensive experiments and demonstrate that our method achieves comparable or better performance than advanced multimodal models, especially showing higher resource efficiency, flexibility, and explainability.
2 Related Work
Harmful Meme Detection
@@ 2 Related Work
aims to classify memes between harmful and harmfless. Contrastive learning-based models e.g., CLIP (Radford et al., 2021) and frameworks have been extensively explored to train multimodal vision language models to fuse cross-modal features and learn unified meme representations in embedding space for classification Mei et al. (2025); Yang et al. (2024); Su et al. (2024); Burbi et al. (2023); Tzelepi and Mezaris (2025); Kumar and Nandakumar (2022). The others convert the original multimodal task to a unimodal text classification setting Ji et al. (2024); Suryawanshi et al. (2023); Cao et al. (2023b) by combining OCR text in memes along with meme image captions generated by pre-trained vision language models e.g., BLIP-2 Li et al. (2023) to train text classifiers based on pure language models such as RoBERTa Liu et al. (2019), FLAN-T5 Chung et al. (2024). To introduce explainability while ensuring classification accuracy, some proposed to explicitly prompt capable LMMs Hee and Lee (2025); Lin et al. (2024a) or LLMsKumari et al. (2024) to perform reasoning before training small MLP or LLM-based classifiers over the generated textual interpretations or rationales with or without other multimodal features. However, existing literature on zero- or few-shot harmful meme detection remains limited. Huang et al. (2024) leverages LLaVA-34B Liu et al. (2023b) and GPT-4o Hurst et al. (2024) as multimodal agents for retrieval-augmented zero-shot harmful meme classification based on a small portion of annotated memes. Despite the potential capabilities seen in large-scale LMMs under such low-resource setting, smaller LMMs with much less parameters yet struggle in zero-shot harmful meme detection Rizwan et al. (2024); Abdullakutty et al. (2024).
LLMs with CoT Reasoning
@@
Thanks to the extensive pretraining on universal knowledge, Large Language Models (LLMs) such as GPT-4, Llama etc., have demonstrated remarkable zero-shot inference abilities over a wide range of reasoning tasks Kojima et al. (2022); Wei et al. (2021) from common-sense, causality and logical reasoning Pan et al. (2023); Qi et al. (2023); Pan et al. (2024); Wu et al. (2024e, f); Wu (2025) to numerical, symbolic and metathetical reasoning Gruver et al. (2023); Miao et al. (2023). The Chain-of-Thought (CoT) approach Wei et al. (2022) improves inference performance by prompting LLMs to explicitly generate intermediate reasoning steps before arriving at a final answer. Recent studies have shown the effectiveness of combining CoT prompting with supervised fine-tuning for meme metaphor detection Xu et al. (2024) and misogynistic meme detection Kumari et al. (2024). Also, Lin et al., 2024b show that some LMMs such as GPT-4V and CogVLM, can benefit from zero-shot CoT prompts for harmful meme classification. However, the capability of unimodal LLMs to perform harmful meme classification under zero-shot CoT prompting remains largely unexplored.
3 Methodology
Since human perceptions and definitions of harmful content (including content that is considered “hateful” or “offensive”, etc.) are often subjective and vary widely with sociocultural contexts, the criteria for detecting harmful memes can differ across platforms and regions and may even evolve over time. Consequently, in real-world applications, it is both natural and reasonable for practitioners to define specific criteria for harmful content they intend to flag and block. This means that the decision boundary between harmful and harmless memes can often be described in natural language. In fact, such criteria often already exist during the data annotation process Pramanick et al. (2021a, b), guiding human annotators in determining whether a meme should be considered harmful. As a result, effective harmful meme detection requires not only commonsense knowledge of widely accepted values, social norms, moral standards, and the historical, cultural, and sociopolitical context of human society but also an understanding of additional context-specific criteria.
General-purpose LLMs, pre-trained on large-scale web corpora and further aligned with human preferences via reinforcement learning, inherently possess extensive world knowledge and fundamental reasoning abilities that enable them to perform well on a wide range of reasoning tasks. Building on this, we hypothesize that even small-scale LLMs internalize sufficient general knowledge and reasoning skills to understand memes. When provided with explicit classification criteria as guidelines, these models can function like human annotators, detecting harmful memes by relying on their inherent capabilities without the need for resource-intensive fine-tuning.
@@ 3 Methodology
With this motivation, we propose U-CoT+, a unimodal guided CoT-based framework that empowers LLMs to perform harmful meme detection via zero-shot inference. Our approach converts the default multimodal detection setup into a unimodal, text-based commonsense reasoning task, allowing LLMs to process textual meme descriptions alongside human-crafted guidelines under zero-shot CoT prompting. This framework comprises two key components: (1) a High-Fidelity Meme2Text pipeline and (2) Unimodal Guided CoT Prompting, as demonstrated in Figure 1.
3.1 High-Fidelity Meme2Text
Harmful memes often target people based on their identity. However, most open-source LMMs tend to produce identity-neutral descriptions when directly prompted to describe images Ji et al. (2024), using ambiguous references such as “a person” or “some people”. This behavior likely results from preference alignment focused on safety concerns. While this may help avoid bias in general applications, it can be problematic for harmful content detection, where identity characteristics of human subjects often provide essential clues.
@@ 3.1 High-Fidelity Meme2Text
In fact, existing harmful meme datasets cover only a limited range of identity characteristics. We identify a few commonly targeted and vulnerable aspects including race, gender, disability, physical appearance, and celebrity status etc., that are critical for detecting harmfulness. To acquire these details, we decompose the task of meme description generation into discrete steps of visual information extraction through basic visual question answering (VQA), which does not necessarily require the underlying models to be highly capable. By repeatedly querying a small LMM (such as Llava1.6-7B), each time focusing on a specific aspect, we can gather all the key visual details of a meme without resorting to more resource-intensive LMMs, while still ensuring comprehensive information coverage necessary for harmful content detection. (See Section C.1 for details.) In this way, we effectively decouple meme content understanding from meme classification reasoning. This helps reduce task complexity by avoiding direct reasoning over complicated, entangled visual elements, which can be prone to hallucination. The proposed high-fidelity meme2text pipeline is demonstrated in Figure 14.
Specifically, for each meme $M$, we obtain a basic low-fidelity description $D_{l}$ through a general “Describe” prompt $P^{D}$ such as “What is shown in the meme?”:
$D_{l}=\mathbf{LMM}(M,P^{D})$ (1)
We then prompt LMMs to identify whether there is any human subject present in the image. For memes that do portray humans, LMMs are futher prompted to provide identify-related information $I_{i}$, where $i$ denotes one of the specific characteristics:
$I_{i}=\mathbf{LMM}(M,P^{I}_{i})$ (2)
To remove verbose expressions and warnings in the outputs of LMMs, rather than merely combining the information, we further leverage unimodal LLMs to integrate all LMM outputs into a coherent and unified description of the meme content. {align} D_h = LLM(D_l, P^I_r, P^I_g, P^I_d, P^I_a, P^I_c) This informative high-fidelity description $D_{h}$ then serves as input for the subsequent text-based classification.
3.2 Unimodal Guided CoT Prompting
3.2.1 Guideline Crafting
The knowledge required for detecting harmful memes is primarily general commonsense. Unlike specialized scientific knowledge that demands expert insights to ensure accuracy, commonsense knowledge is widely accessible and does not require significant cost to acquire or verify. Consequently, crafting harmful meme detection guidelines manually is both intuitive and cost-effective, eliminating the need for heavy reliance on experts to identify patterns of harmful memes in existing datasets. Grounded in this empirical commonsense, we first collect and analyze the definitions, explanations, and examples of harmful memes provided by the authors in each dataset paper, along with their reported data annotation guidelines as initial references. From this analysis, we identify several types of guidelines that are shared across sociocultural contexts and can support LLMs in reasoning more effectively when detecting harmfulness in memes.
Implicitness
Memes containing derogatory language, abusive slurs, offensive label-calling, direct personal attacks or indication of hate are explicitly harmful, making them easy to flag by LLMs that have been pre-trained to warn about the presence of profanity or swearing. However, as highlighted in Huang et al. (2024), many harmful memes included in existing datasets can be implicit, with their harmful messages wrapped under the deceptive disguise of benign visuals and neutral language. Therefore, this guideline requires LLMs to carefully reflect on a meme’s implication to check whether it is deliberately crafted to evoke negative contextual interpretations e.g., associations to harmful stereotypes, sensitive controversies in societal, cultural or political topics etc. , that are actually meant to be malicious.
Tone & Intent
We observe that both multi- and uni-modal LLMs tend to interpret an arbitrary meme positively if they are not explicitly reminded. Such behavior is probably the result of human preference alignment that aims to ensure appropriate and safe responses, but is undesirable for harmful meme classification. To this end, we include this guideline to remind LLMs not to presume the nature of a meme’s tone and intent as humorous, playful or light-hearted in the first place, but to interpret from a neutral perspective.
Fine-grained Taxonomy

### v2

# Read as You See: Guiding Unimodal LLMs for Low-Resource Explainable Harmful Meme Detection

@@
Abstract.
Detecting harmful memes is crucial for safeguarding the integrity and harmony of online environments, yet existing detection methods are often resource-intensive, inflexible, and lacking explainability, limiting their applicability in assisting real-world web content moderation. We propose U-CoT+, a resource-efficient framework that prioritizes accessibility, flexibility and transparency in harmful meme detection by fully harnessing the capabilities of lightweight unimodal large language models (LLMs). Instead of directly prompting or fine-tuning large multimodal models (LMMs) as black-box classifiers, we avoid immediate reasoning over complex visual inputs but decouple meme content recognition from meme harmfulness analysis through a high-fidelity meme-to-text pipeline, which collaborates lightweight LMMs and LLMs to convert multimodal memes into natural language descriptions that preserve critical visual information, thus enabling text-only LLMs to “see” memes by “reading”. Grounded in textual inputs, we further guide unimodal LLMs’ reasoning under zero-shot Chain-of-Thoughts (CoT) prompting with targeted, interpretable, context-aware, and easily obtained human-crafted guidelines, thus providing accountable step-by-step rationales, while enabling flexible and efficient adaptation to diverse sociocultural criteria of harmfulness. Extensive experiments on seven benchmark datasets show that U-CoT+ achieves performance comparable to resource-intensive baselines, highlighting its effectiveness and potential as a scalable, explainable, and low-resource solution to support harmful meme detection [^1] Codes and data are available at: https://github.com/panFJCharlotte98/HMC.
Harmful meme detection; Low-resource; Explainability; LLMs; LMMs
journalyear: 2026conference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates
1. Introduction
@@
As a popular communication medium on social media platforms (Nguyen and Ng, 2024; Joshi et al., 2024), memes convey rich information in a visually engaging manner by combining images with overlaid text (Dancygier and Vandelanotte, 2017). Despite their often amusing nature, memes can be exploited to disseminate harmful content under the guise of humor, such as hate speech (Kiela et al., 2020; Davidson et al., 2019), toxic misinformation (Lu et al., 2024; Kennedy et al., 2022), and inflammatory opinions (Liu et al., 2022; Edwards and Gribbon, 2013). Harmful memes [^2] Disclaimer: This paper contains demonstration content that may be disturbing. pose a threat to the safety and harmony of online environments (Pandiani et al., 2024; Sharma et al., 2022), thus motivating automated harmful meme detection to support content moderation for online platforms.
@@
However, existing methods, most of which follow a typical supervised fine-tuning (SFT) paradigm, face three critical limitations: (i) Poor Resource Efficiency. On one hand, supervised approaches tend to be data-inefficient as they require large volumes of labeled training data to perform well (Burbi et al., 2023; Shah et al., 2024; Kumar and Nandakumar, 2022; Pramanick et al., 2021b). However, the nuanced, implicit and context-specific nature of memes (Chakravarthi et al., 2025; Baruah et al., 2020) makes it difficult for human annotators to maintain consistent labeling without being limited by individual knowledge and subjectivity (Gillespie, 2020). As such, collecting and annotating memes are often time-consuming, costly, and prone to bias. On the other hand, existing low-resource methods (Huang et al., 2024; Ji et al., 2024; Lin et al., 2024a) often rely on proprietary Large Multimodal Models (LMMs) e.g., GPT-4 or heavily scaled models with massive parameter counts, incurring substantial financial and computational expenses. Taken together, these methods pose significant accessibility challenges for practitioners operating under low-resource constraints. (ii) Limited Scalability and Flexibility. The classification criteria distinguishing harmful and harmless memes often vary across sociocultural contexts (Zannettou et al., 2018; Sharma et al., 2022) and differ across platforms, regions, and over time (Bonagiri et al., 2025; Singhal et al., 2023; Pater et al., 2016; Gorwa et al., 2020), which entails the use of specialized policies for different scenarios in order to robustly define harmful memes (Thomas et al., 2025). However, supervised classifiers are not easily scalable. Once trained, they become difficult to adapt to new out-of-domain criteria, meme or harmfulness types without undergoing costly data collection and retraining (Mei et al., 2025; Cao et al., 2024). (iii) Lack of Explainability. Real-world content moderation practices combine model-generated decisions to assist human verdicts (Bonagiri et al., 2025; Facebook, 2024). However, most supervised methods operate as black boxes, offering no human-interpretable reasoning behind their predictions (Burkart and Huber, 2021; Yang et al., 2024; Burbi et al., 2023). This lack of transparency in decision-making undermines trust in human-machine collaboration and deployment reliability in sensitive or regulated scenarios (Sasikala and Sachan, 2024), where explicit rationales are essential to support subsequent human reviews and feedback.
@@
To address these limitations, in this paper, we explore the feasibility of harnessing lightweight Large Language Models (LLMs) for low-resource harmful meme detection. Specifically, we propose U-CoT+, a unimodal guided Chain-of-Thoughts (CoT) (Wei et al., 2022) prompting framework, which converts the default multimodal setup back to unimodal text-only inference, enabling LLMs to perform zero-shot harmful meme detection solely from textual input. To this end, we first propose a High-Fidelity Meme2Text pipeline. Instead of directly prompting LMMs to perform classification, we query lightweight LMMs multiple times to extract critical visual cues (Cao et al., 2023a; Ji et al., 2024) for sensitive, identity-related attributes of human subjects (e.g., race, gender). We then resort to a unimodal LLM to integrate the LMM outputs into a unified, informative description which it subsequently uses for detection. This design effectively decouples content recognition from harmfulness analysis, reducing task complexity by avoiding immediate reasoning over complicated visual elements.
Furthermore, we introduce the Guided CoT approach. Rather than relying solely on LLMs’ internalized knowledge, we guide them with explicit, human-crafted guidelines that specify fine-grained, targeted criteria for distinguishing harmful from harmless memes. LLMs reason step by step over textual meme descriptions in conjunction with the guidelines, promoting consistent, interpretable, and aligned classifications. This also allows easy and effective adaptation to different harmfulness criteria across diverse sociocultural contexts with only minor resources and human effort.
Through extensive experiments on benchmark datasets, we show that our proposed zero-shot harmful meme detection framework effectively narrows the performance gap to state-of-the-art full-shot supervised fine-tuned baselines. Notably, with guided CoT prompting, lightweight LLMs (such as Mistral-12B and Qwen2.5-14B) perform competitively on zero-shot inference, matching and even outperforming GPT-4o(-mini) and other low-resource baselines. We summarize our main contributions as three-fold:
@@
  • We propose a harmful meme detection framework that effectively leverages the inherent image understanding abilities, commonsense knowledge and reasoning abilities of lightweight open-source LMMs and LLMs to enable explainable meme classification under low-resource settings. By performing zero-shot inference on high-fidelity meme descriptions under guided CoT prompting, our framework can easily adapt to different harmfulness evaluation criteria across diverse sociocultural contexts.
@@
  • We contribute a set of interpretable and targeted guidelines for harmful meme detection across various contexts. We demonstrate that these high-quality yet simple, human-crafted guidelines, designed to capture diverse forms and features of harmfulness, can effectively enhance the zero-shot performance of LLMs in detecting harmful memes.
  • We benchmark the zero-shot harmful meme detection performance of representative lightweight unimodal LLMs. We show that our method fully leverages their inherent capabilities, achieving performance comparable and even superior to that of resource-intensive multimodal methods. This highlights an accessible yet effective solution for multimodal tasks in low-resource scenarios.
Figure 1. Illustration of previous methods and our U-CoT+. Previous methods follow either fully supervised or low-resource settings, fine-tuning LMMs/PLMs with labeled data or prompting advanced LMMs (e.g., GPT-4) with/without few-shot examples or retrieval-augmented mechanisms. They do not necessarily guarantee predictions that include explicit reasoning. U-CoT+ employs a High-fidelity Meme2Text pipeline to convert the multimodal harmful meme detection task into a unimodal, text-only setting, and further enhances LLMs’ reasoning through Unimodal Guided CoT Prompting. An example output given by Qwen2.5-14B under U-CoT+ is presented in “Step-by-step Reasoning”. The meme is only for demonstration purposes.
2. Related Work
Harmful Meme Detection.
@@ 2. Related Work
This task seeks to classify online memes as harmful or harmless. A major line of works explore contrastive learning models and frameworks (e.g., based on CLIP (Radford et al., 2021)) to train vision language models as multimodal classifiers that fuse cross-modal features and learn unified meme representations in embedding space for image classification  (Mei et al., 2025; Yang et al., 2024; Su et al., 2024; Burbi et al., 2023; Tzelepi and Mezaris, 2025; Kumar and Nandakumar, 2022). Others generate image captions and fine-tune binary text classifiers based on pre-trained language models (Ji et al., 2024; Suryawanshi et al., 2023; Cao et al., 2023b). Some injects explainability in detection by prompting advanced LMMs (Hee and Lee, 2025; Lin et al., 2024a) or LLMs (Kumari et al., 2024) to generate reasoning outputs for supervised training. However, existing work on low-resource harmful meme detection remains limited. The use of advanced multimodal models such as LLaVA-34B and GPT-4o as training-free detection agents is explored in (Huang et al., 2024). Yet, small-scale LMMs struggle with zero-shot detection (Rizwan et al., 2024; Abdullakutty et al., 2024).
@@
LLMs with CoT Reasoning. LLMs extensively pretrained on universal knowledge have demonstrated remarkable zero-shot inference abilities over manifold reasoning tasks (Kojima et al., 2022; Wei et al., 2021; Qi et al., 2023; Wu et al., 2024c; Pan et al., 2024; Wu et al., 2024e, a, b, d; Wu, 2025; Wu et al., 2025). The Chain-of-Thought (CoT) method (Wei et al., 2022) aims to improve inference performance by prompting LLMs to explicitly generate intermediate reasoning steps before arriving at a final verdict. Recent studies have shown the effectiveness of CoT-enhanced supervised fine-tuning and LMMs (e.g., GPT-4V) for meme-centered tasks (Xu et al., 2024; Kumari et al., 2024; Lin et al., 2024b). However, whether and how this applies to lightweight unimodal LLMs for harmful meme detection remains largely unexplored.
3. Methodology
The criteria for defining harmful content vary widely with sociocultural contexts and differ across platforms, regions and even over time. The use of diverse specialized policies in real-world content moderation highlights the need for context/domain-specific guidelines. By nature, harmful meme detection is a linguistic reasoning task that relies on the understanding of: (i) Nuances in language use, including metaphors, sarcasm, humor and other rhetorical devices; (ii) Commonsense knowledgesuch as social norms, moral values, sociopolitical factors and historical, cultural context; (iii) Explicit criteriathat define the decision boundary between harmful and benign content. Unimodal text-only LLMs, pre-trained on large-scale web corpora and further aligned with human preferences, inherently possess extensive world knowledge and fundamental reasoning abilities. Trained with refined objectives specifically optimized for general-purpose reasoning and instruction following, they demonstrate more reliable performance on language-only tasks than multimodal counterparts of comparable size. Their availability at various scales also offers flexible options for low-resource deployment.
@@ 3. Methodology
Therefore, we propose U-CoT+, a unimodal guided CoT framework that aims to harness the inherent abilities of lightweight LLMs in zero-shot inference for low-resource harmful meme detection. Adopting an empirically grounded, multi-stage LMM+LLM collaborative workflow (Jha et al., 2024; Adak et al., 2025; Agarwal et al., 2024), our approach effectively converts the default multimodal classification task into a unimodal, text-based linguistic reasoning setup, allowing LLMs to process memes based on textual descriptions and analyze meme harmfulness guided by human-crafted guidelines. The framework comprises two key components: (1) a High-Fidelity Meme2Text pipeline and (2) Unimodal Guided CoT Prompting, as demonstrated in Figure 1.
3.1. High-Fidelity Meme2Text
@@
Typical harmful memes use brief, salient visual summaries to attack people’s identities. Interpreting their punchlines does not require exhaustive, irrelevant visual details, but rests on a clear understanding of what or who is being targeted. Commonly found vulnerable identity-related attributes critical for harmfulness detection include race, gender, disability, physical appearance, and celebrityetc.  However, safety-aligned LMMs, due to debiasing considerations in general-purpose applications, tend to omit socially and culturally sensitive visual cues but use ambiguous, identity-neutral references (e.g., “a person” or “some people”) when describing images with human subjects (Ji et al., 2024), which poses challenges for harmful content detection. To retain such essential content, we decompose meme description generation into individual steps of visual information extraction via atomic question answering (VQA), which does not necessarily require highly capable models. Therefore, by repeatedly querying a lightweight LMM (such as LLaVA1.6-7B), each time focusing on a specific attribute, we can gather these key information without resorting to more resource-intensive LMMs. In this way, we effectively decouple meme content recognition from meme classification reasoning. This helps reduce task complexity by avoiding immediate reasoning over complicated, entangled visual elements, which is prone to hallucination.
Specifically, for each meme $M$, LMMs are given a general “Describe” prompt $P^{D}$ such as “What is shown in the meme?” to generate a basic low-fidelity description: $D_{l}=\mathbf{LMM}(M,P^{D})$, and then identify the presence of any human subject. For memes showing humans, LMMs are further prompted to provide identify-related information $I_{i}=\mathbf{LMM}(M,P^{I}_{i})$, where $i$ denotes different attributes. With all visual cues ready, we further leverage unimodal LLMs to integrate all LMM outputs into a unified, compact, and coherent high-fidelity description of the meme content: $D_{h}=\mathbf{LLM}(D_{l},I_{r},I_{g},I_{d},I_{a},I_{c})$, which serves as input for the subsequent text-based classification.
3.2. Unimodal Guided CoT Prompting
3.2.1. Guideline Crafting
The subjective and context-dependent nature of harmful meme definitions makes it unlikely to eliminate human involvement in guideline crafting process. However, harmful meme detection largely relies on widely accessible commonsense knowledge rather than specialized expert insights. Accordingly, we propose an empirical guideline crafting procedure that incurs only minor human effort: (i) We examine the available harmful meme definitions and data annotation guidelines in existing datasets for the patterns of harmfulness in representative memes; (ii) We summarize the observed patterns and underlying criteria into coherent natural language guidelines consisting of multiple rules ordered in a logically smooth way (e.g., from general to specific, from features to examples, etc.); (iii) We refine guidelines by probing their clarity and effectiveness on sample memes. This procedure yields the following set of principles for guideline formulation that are reproducible and transferable across different sociocultural contexts:
  • Implicitness Harmful memes can be implicit (Huang et al., 2024). This principle requires LLMs to reflect on the underlying implications of a meme to examine whether it is deliberately crafted to evoke negative contextual interpretations (e.g., associations to harmful stereotypes, sensitive controversies in societal, cultural or political practices etc.) while wrapping harmful messages beneath deceptively benign visuals and neutral language.
  • Tone & Intent LLMs tend to interpret memes charitably due to alignment with human preferences for safe responses. To mitigate such tendency, this principle explicitly instructs LLMs to avoid assuming a playful, humorous or light-hearted tone and intent, and instead interpret from a neutral perspective.